OpenStudy (anonymous):

jenn has $150 saved for lunch money. each day she takes out $2.50 for lunch write the equation for this function???D:

OpenStudy (luigi0210):

What we know so far is that the equation is linear, correct?

OpenStudy (anonymous):

im sorry replace the word equation with formula the section im in is called arithmetic and geometric sequencing if that helps

OpenStudy (luigi0210):

we could still y=mx+b if I'm not mistaken

OpenStudy (anonymous):

sounds about right

OpenStudy (luigi0210):

Okay, so far we know that we have two numbers. The total of 150 and 2.50, now we need to figure out which one gets the variable x

OpenStudy (anonymous):

wouldnt 2.50 get the x?

OpenStudy (luigi0210):

Yes it would, and instead of it being y=mx+b it's gonna be y=b-mx because we are taking out money not adding some.

OpenStudy (mathstudent55):

Let m = money Jenn has She starts with m = 150 before spending any money on lunches On day 1, she takes out $2.50, so she ends up with m = 150 - 2.50 On day 2, she starts with m = 150 - 2.50 = 147.50, and she spends another 2.50, so she ends up with m = 147.50 - 2.50, but this is also m = 150 - 2.50 - 2.50 or m = 150 - 2*2.50, She took out 2 times 2.50, one times 2.50 on day one and one times 2.50 on day 2. On day 3, she ends up with 150 - 2.50 - 2.50 - 2.50 or m = 150 - 3*2.50 Let d = number of the day On any day d, she will end up with m = 150 - d*2.50 after paying for lunch. The formula then is m = 150 - 2.5d Where m = the money she will have after paying for lunch on day d. Now that you have the formula, a question you can ask is how many days will it take for her to spend all her money. That means after spending 2.50 in the last day, she will have 0 dollars left. That is when m = 0. Solve 0 = 150 - 2.5d for d, to find out.

OpenStudy (luigi0210):

... Yea pretty much what mathstudent explained :l